Calculate the molar mass of cuso4.5h2o
1
Expert's answer
2020-12-18T08:47:20-0500

Mr(CuSO4 * 5H2O)= Mr(CuSO4) + 5Mr(H2O)= 64 +32 + 64 + 5*(2 + 16)= 250


Answer: 250 g/mole